A sample of B2H6 gas in a 35.0-mL container is at a pressure of 450. mm Hg and a temperature of 571 °C. If the gas is allowed to cool at constant volume until the pressure is 354 mm Hg, what is the new temperature in °C?


1
Expert's answer
2021-03-15T09:28:47-0400

"\frac{P_1}{T_1}=\frac{P_2}{T_2}"


="\frac{450}{844}=\frac{354}{T_2}"


="T_2=\frac{844×354}{450}" =663.95K


"T_2=390.95^{o}C"